### v1

# Optical Properties, Quasinormal Modes and Greybody factors of deformed AdS-Schwarzschild black holes

Abstract
We investigate the temperature, photon and shadow radii, quasinormal modes (QNMs), greybody factors, and emission rates of deformed AdS black holes, focusing on the effects of the deformation parameter $\alpha$ and control parameter $\beta$. Increasing $\alpha$ enhances the oscillation frequency and damping rate of gravitational waves, while $\beta$ shows non-linear behaviour. Electromagnetic perturbations exhibit similar trends, though with lower frequencies and damping rates. Greybody factors are mainly influenced by multipole moment $l$ and $\alpha$, with $\beta$ having a more subtle effect. These findings provide insights into black hole dynamics, mergers, and gravitational wave emissions.
WKB method; Black hole; Quasinormal modes, Greybody factors
pacs:
04.30.Tv, 04.50.Kd, 97.60.Lf, 04.70.-s
I Introduction
The study of black holes has captivated scientists for decades due to their extreme gravitational properties, but their importance extends beyond gravity. Black holes are now recognized as crucial objects for understanding thermodynamic principles in the most extreme conditions of the universe. Black hole thermodynamics, a field that merges the principles of classical thermodynamics with general relativity, quantum mechanics, and statistical mechanics, has emerged as a fascinating and vital area of research. It not only deepens our understanding of black hole behaviour but also serves as a potential gateway to formulating a comprehensive theory of quantum gravity.
Black hole thermodynamics began with a groundbreaking insight by J. D. Bekenstein, who, in the early 1970s, proposed that black holes should possess entropy. His work established a direct relationship between the area of a black hole’s event horizon and its entropy, indicating that the surface area of the horizon could be a measure of the disorder or information content of a black hole Bekenstein:1972tm . This idea transformed the way scientists viewed black holes, not merely as objects that consume matter and energy, but as thermodynamic entities that obey laws analogous to those of classical thermodynamics.
Building upon Bekenstein’s work, Stephen Hawking made a revolutionary discovery in 1974 when he demonstrated that black holes can emit radiation, now famously known as Hawking radiation. This radiation arises from quantum mechanical effects near the event horizon, where particle-antiparticle pairs are created, and one of the particles escapes while the other falls into the black hole. Over time, this process causes the black hole to lose mass and eventually evaporate Hawking:1975vcx . Hawking’s discovery challenged the traditional notion of black holes as objects from which nothing can escape and introduced a connection between black hole thermodynamics and quantum mechanics. This discovery also implied that black holes possess a finite temperature, related to their mass and surface area, adding to the richness of their thermodynamic properties Wald:1999vt .
The relationship between the area of a black hole’s event horizon and its entropy, formulated by Bekenstein and Hawking, is now considered a cornerstone of black hole thermodynamics. Their collaboration helped define the four laws of black hole thermodynamics, which closely parallel the classical laws of thermodynamics. These laws govern the behaviour of black hole systems in terms of temperature, entropy, and energy conservation Bekenstein:1973ur ; Wald:1999vt ; Bardeen:1973gs ; Georgescu:2021dyo . Specifically, the second law of black hole thermodynamics, which states that the total entropy of a black hole cannot decrease, is analogous to the second law of thermodynamics, which dictates that the total entropy of a closed system never decreases.
As the field progressed through the 1980s and 1990s, black hole thermodynamics became more intricately connected with quantum mechanics. One of the most significant developments during this period was the introduction of the holographic principle, which postulates that all the information contained within a volume of space can be represented on the boundary of that space. This idea suggests that the physics inside a black hole, or even a more general region of spacetime, can be understood in terms of information encoded on its boundary, hinting at deeper connections between gravity and quantum field theory Wald:1999vt . This concept found concrete application in the AdS/CFT correspondence, a powerful tool in theoretical physics that relates a gravitational theory in anti-de Sitter (AdS) space to a conformal field theory (CFT) on its boundary. Such correspondences have proven invaluable in studying black hole thermodynamics and quantum gravity.
In recent years, research has continued to explore black hole thermodynamics across a variety of theoretical frameworks, including modified theories of gravity such as string theory and loop quantum gravity Ashtekar:2004eh ; Dvali:2011aa . These modified theories often lead to generalizations or deformations of the traditional thermodynamic laws that apply to black holes. For example, when incorporating the generalized uncertainty principle, a modification of the Heisenberg uncertainty principle to account for quantum gravitational effects, the entropy and temperature of black holes are modified, leading to potentially observable consequences Amelino-Camelia:2000stu ; Ali:2009zq ; Yang:2023nnk . These generalized uncertainty principles are particularly relevant in the context of small black holes or black holes that form in high-energy regimes, such as those potentially produced in particle accelerators.
Furthermore, external factors such as electromagnetic fields or the presence of a cosmological constant can significantly alter the thermodynamic behaviour of black holes. For instance, the presence of a cosmological constant, which leads to anti-de Sitter or de Sitter space solutions, introduces new thermodynamic properties and phase transitions in black holes. Such phase transitions, which are akin to those found in classical thermodynamics (such as the transition between solid, liquid, and gas phases), are observed in various black hole solutions, including charged, rotating, and higher-dimensional black holes Kubiznak:2012wp ; Hendi:2012um ; Teitelboim:1985dp ; Gunasekaran:2012dq ; Cvetic:1999ne ; Chamblin:1999tk . In these systems, changes in parameters like charge, angular momentum, or cosmological constant can lead to critical phenomena and phase transitions, which have been a topic of intense study in recent years.
For example, research on rotating Kerr-AdS black holes using holographic techniques has provided insights into their phase transitions and thermodynamic properties Gong:2023ywu . Similarly, accelerating black holes and those in lower-dimensional spacetimes offer simplified contexts for studying thermodynamic stability and transitions, contributing to the broader understanding of black hole microstructures Sokoliuk:2023pby ; Singh:2023hit ; Blagojevic:2023kqj ; Xiao:2023lap ; Wu:2022whe . Such studies are crucial for understanding the nature of black hole entropy and the information paradox, one of the most significant unresolved issues in theoretical physics.
Phase transitions and critical phenomena have been extensively investigated in black hole thermodynamics Gunasekaran:2012dq ; Cvetic:1999ne ; Chamblin:1999tk . Various types of black holes in different modified gravity theories have also been found to exhibit a wide range of phase transitions Wei:2012ui ; Altamirano:2014tva .
Through holographic techniques, rotating Kerr-AdS black holes have been analyzed for their phase transitions and thermodynamic properties Gong:2023ywu , while three-dimensional accelerating black holes provide a more simplified context for such studies Tian:2023ine . The authors of Ref. Sokoliuk:2023pby examined AdS black holes, focusing on their microstructures and stability, contributing to insights related to the AdS/CFT correspondence. In Ref. Singh:2023hit , thermodynamic curvature transitions across different ensembles in charged AdS2 black holes were explored. Charged black holes with scalar fields challenge the no-hair theorem, affecting their entropy, temperature, and phase transitions, as discussed in Ref. Blagojevic:2023kqj . Furthermore, the Iyer-Wald formalism has been used to derive thermodynamic laws for AdS-Kerr black holes, addressing controversies in the field, as studied in Ref. Xiao:2023lap . The impact of topological properties on black hole thermodynamic stability has also been considered Wu:2022whe ; Wei:2021vdx ; Rizwan:2023ivp . Critical points of charged Gauss-Bonnet black holes in AdS were classified in Ref. Yerra:2022alz . Additional studies include the tunneling of charged bosons from noncommutative charged black holes, from which the rate and Hawking temperature were derived Ovgun:2015box , the formulation of thermodynamics and computation of dual stress-energy tensors for accelerating black holes in AdS Anabalon , and the generalization of non-relativistic black hole solutions to include electric charge, exploring their properties Jusufi .
In summary, black hole thermodynamics not only enhances our understanding of the fundamental properties of black holes but also plays a central role in the quest for a theory of quantum gravity. Through the study of black hole entropy, temperature, and phase transitions, physicists are uncovering profound connections between gravity, quantum mechanics, and thermodynamics. As research continues to evolve, black hole thermodynamics remains a fertile ground for exploring some of the deepest questions in theoretical physics.
In addition to the thermodynamic properties of black holes, their quasinormal modes (QNMs) and greybody factors are critical for understanding the dynamics and observational signatures of these enigmatic objects. Quasinormal modes represent the characteristic oscillations that black holes undergo when they are perturbed Konoplya:2003ii ; Lambiase:2024lvo ; Konoplya:2023ppx ; Konoplya:2023aph ; Konoplya:2022hll ; Konoplya:2019hlu . These oscillations are damped over time due to the emission of gravitational waves, and the frequencies and damping rates of the QNMs are determined by the black hole’s mass, charge, and angular momentum. The study of QNMs provides invaluable insights into the stability of black holes, as well as their interactions with external perturbations. Furthermore, QNMs are directly related to the ringdown phase of gravitational wave signals detected by observatories such as LIGO and Virgo. By analyzing the QNMs from observed gravitational waves, scientists can infer the parameters of the black hole, offering a unique observational tool for testing general relativity in the strong-field regime and exploring potential deviations from it in modified theories of gravity.
Closely related to QNMs are greybody factors, which describe the modification of Hawking radiation as it escapes from the black hole’s gravitational well. In the absence of any scattering, Hawking radiation would have a perfect blackbody spectrum, but greybody factors account for the fact that some of the radiation is scattered back into the black hole by the spacetime curvature. This scattering alters the spectrum of the emitted radiation and provides important corrections to the pure Hawking radiation spectrum. The study of greybody factors is crucial for understanding the energy emission rates of black holes and for exploring potential observational signatures of black holes in the universe. Together with QNMs, greybody factors enrich our understanding of black hole thermodynamics and play a key role in connecting theoretical predictions with astrophysical observations, particularly in the era of gravitational wave astronomy.
In this paper, we adopt the metric signature (,+,+,+)(-,+,+,+)( - , + , + , + ) and use geometrical units where $c=G=\hbar=1$. After this brief introduction, the paper is structured as follows:
In Section II, we provide a concise overview of the deformed black hole solution. Section III examines the Hawking temperature of the black hole and explores the remnant radius and mass. In Section IV, we analyze the photon sphere radius based on the motion of light and derive the black hole’s shadow radius. Section V investigates the quasinormal modes (QNMs) of the black hole spacetime, while Section VI focuses on the greybody factors. The emission rate of the black hole is discussed in Section VII. Finally, in Section VIII, we conclude with a brief summary and discussion.
II Deformed Black hole solution
In this section, we explore the construction of a deformed Anti-de Sitter (AdS) black hole solution within a modified gravity framework. The action governing the four-dimensional spacetime is given by Khosravipoor
$A=\int d^{4}x\,\sqrt{-g}\left(\frac{R-2\Lambda}{2\kappa}+\mathcal{L}_{m}+\mathcal{L}_{\rm X}\right),$ (1)
where $R$ is the Ricci scalar, $\Lambda$ is the cosmological constant, $\kappa=8\pi G/c^{4}$, and $\mathcal{L}_{m}$ and $\mathcal{L}_{\rm X}$ represent the matter and beyond-General Relativity (GR) Lagrangians, respectively. The term $\mathcal{L}_{\rm X}$ encompasses contributions from extra fields such as scalar, vector, or tensor fields, or extensions of gravity theories that modify the standard Einstein-Hilbert action. When $\mathcal{L}_{\rm X}=0$, the resulting solution corresponds to the familiar AdS-Schwarzschild metric. However, the presence of $\mathcal{L}_{\rm X}$ leads to a deformed black hole solution with interesting physical implications.
The Einstein field equations arising from this action can be written as
$G_{\mu\nu}+\Lambda g_{\mu\nu}=\kappa\,T^{(\rm tot)}_{\mu\nu},$ (2)
where $G_{\mu\nu}$ is the Einstein tensor, and the total stress-energy tensor $T^{(\rm tot)}_{\mu\nu}$ includes contributions from both the matter content and the extra fields described by $\mathcal{L}_{\rm X}$, i.e.,
$T^{(\rm tot)}_{\mu\nu}=T^{(\rm m)}_{\mu\nu}+T^{(\rm X)}_{\mu\nu}.$ (3)
To construct the deformed black hole solution, we consider a static, spherically symmetric spacetime with the line element
$ds^{2}=-\mathcal{F}(r)dt^{2}+\dfrac{1}{\mathcal{F}(r)}dr^{2}+r^{2}(d\theta^{2}+\sin^{2}{\theta}\,d\phi^{2}).$ (4)
The deformation in the black hole geometry is introduced via an energy density function $\mathcal{E}(r)$, which is chosen to be proportional to $1/r^{4}$, as follows Khosravipoor :
$\mathcal{E}(r)=\frac{\alpha}{\kappa(\beta+r)^{4}},$ (5)
where $\alpha$ and $\beta$ are deformation parameters. The parameter $\alpha$ controls the strength of the deformation, while $\beta$ ensures the avoidance of central singularities at $r=0$. This choice of energy density function satisfies the necessary energy conditions and asymptotically decays as $r\to\infty$.
Substituting $\mathcal{E}(r)$ into the field equations yields the modified metric function for the deformed AdS black hole Khosravipoor :
$\mathcal{F}(r)=1-\frac{2M}{r}-\frac{\Lambda r^{2}}{3}+\alpha\,\frac{\beta^{2}+3r^{2}+3\beta r}{3r(\beta+r)^{3}}.$ (6)
where $M$ is the mass of the black hole. The asymptotic form of the metric function remains consistent with the AdS-Schwarzschild solution, but the additional term proportional to $\alpha$ introduces deviations that modify the black hole’s properties.
The deformed AdS black hole solution offers a rich framework for exploring deviations from GR in strong gravitational regimes. The modifications introduced by $\alpha$ and $\beta$ affect key aspects of black hole physics, such as quasinormal modes (QNMs), greybody factors, and gravitational wave emission during black hole mergers. The variations in the black hole’s response to perturbations, particularly in the ringdown phase of mergers, provide potential observational signatures that could help distinguish deformed black holes from their standard counterparts.
Moreover, the inclusion of extra fields or modifications to the gravitational sector, as encapsulated by $\mathcal{L}_{\rm X}$, aligns with ongoing efforts to reconcile GR with quantum gravity and other beyond-standard-model physics. By examining the behaviour of the deformed black hole in both classical and quantum contexts, we can gain deeper insights into the nature of gravity and the possible extensions of Einstein’s theory.
In the following sections, we will investigate the quasinormal modes and greybody factors associated with this deformed black hole solution and explore their observational consequences.
III Temperature
At horizon $\mathcal{F}(r)=0\big{|}_{r=r_{h}}$, thus one can find the mass of the black hole as a function of the horizon radius.
$M(r_{h})=-\frac{1}{2}r_{h}\left(-\frac{\alpha\left(\beta^{2}+3r_{h}^{2}+3\beta r_{h}\right)}{3r_{h}(\beta+r_{h})^{3}}+\frac{\Lambda r_{h}^{2}}{3}-1\right).$ (7)
Hawking temperature is obtained from $\frac{1}{4\pi}\partial\mathcal{F}(r)/\partial r|_{r=r_{h}}$. For the metric of the form 6, by substituting mass from Eq. (7), the temperature would be found as
$T_{H}=\frac{\beta^{4}+r^{4}+4\beta r^{3}-\alpha r^{2}+6\beta^{2}r^{2}-\Lambda r^{2}(\beta+r)^{4}+4\beta^{3}r}{4\pi r(\beta+r)^{4}}$ (8)
The temperature curve is illustrated in Fig. 1.
Refer to caption      Refer to caption
Figure 1: Variation of the Hawking temperature with radial distance for $\beta=0.2,\,\Lambda=-0.002$ (right panel) and $\alpha=0.2,\,\Lambda=-0.002$ (left panel) respectively.
We are interested in cases that temperature curve crosses $r_{h}$ axis in a positive value of $r_{h}$, this point is called remnant radius and is calculated as
$T_{H}\Big{|}_{r=r_{rem}}=\frac{\beta^{4}+r_{rem}^{4}+4\beta r_{rem}^{3}-\alpha r_{rem}^{2}+6\beta^{2}r_{rem}^{2}-\Lambda r_{rem}^{2}(\beta+r_{rem})^{4}+4\beta^{3}r_{rem}}{4\pi r_{rem}(\beta+r_{rem})^{4}}=0$ (9)
Remnant radius as a function of parameter $\alpha$ is demonstrated in Fig. 2
Figure 2: Behaviour of remnant radius as a function of parameter $\alpha$ with $\Lambda=-0.002$.
As is apparent, while with $\beta$, and $\Lambda$ held constant, an increase in the parameter $\alpha$ results in an enlargement of the remnant radius, with $\alpha$, and $\Lambda$ held constant, an increase in the parameter $\beta$ leads to a decrease in the value of the remnant radius.
By substituting remnant mass from Eq. (9), by $r_{h}$ in Eq. (7), remnant mass of the black hole can be find as
$M(r_{rem})=-\frac{1}{2}r_{rem}\left(-\frac{\alpha\left(\beta^{2}+3r_{rem}^{2}+3\beta r_{rem}\right)}{3r_{rem}(\beta+r_{rem})^{3}}+\frac{\Lambda r_{rem}^{2}}{3}-1\right).$ (10)
IV Photon and Shadow Radii
Assuming a massless case, the Euler-Lagrange equation is written as
$\frac{d}{d\tau}(\frac{\partial\mathcal{L}}{\partial\dot{x}^{\mu}})-\frac{\partial\mathcal{L}}{\partial x^{\mu}}=0,$ (11)
where Lagrangian is obtained as Ahmed
$\mathcal{L}=\frac{1}{2}g_{\mu\nu}\dot{x}^{\mu}\dot{x}^{\nu}.$ (12)
metric tensor $g_{\mu\nu}$ can be calculated from Eq. (4). In the equatorial plane where $\theta=\pi/2$, the Lagrangian yields
$2\mathcal{L}=-\mathcal{F}(r)\dot{t}^{2}+\frac{\dot{r}}{\mathcal{F}(r)}+r^{2}\dot{\phi}^{2}$ (13)
Assuming two motion constants as $E=\partial\mathcal{L}/\partial\dot{t}=\mathcal{F}(r)\dot{t}$ and $L=\partial\mathcal{L}/\partial\dot{\phi}=r^{2}\dot{\phi}$, where $E$ referred to the energy and $L$ indicates the angular momentum, for null geodesic Eq. (11) reduces to
$\left(\frac{dr}{d\tau}\right)^{2}+\frac{L^{2}}{r^{2}}\mathcal{F}(r)-E^{2}=0.$ (14)
Thus one can define an effective potential which is formulated as
$V_{eff}(r)=\mathcal{F}(r)\frac{L^{2}}{r^{2}}-E^{2}$ (15)
To ensure a circular orbit, two constraints are applied which are as follow
$V_{Eff}(r)=0,\;\;\partial_{r}V_{Eff}(r)=0,$ (16)
Thus one can find that the following constraints should be satisfied
$E/L$ =±$=\pm\sqrt{\mathcal{F}(r)/r^{2}},$ (17)
$r^{2}\partial_{r}$ $\mathcal{F}(r)-2r\mathcal{F}(r)=0$ (18)
By solving Eq. (18) at $r=r_{ph}$, one can find the photon radius of the black hole.
Shadow radius is given by
$r_{sh}=r_{ph}\frac{\sqrt{f(r_{o})}}{\sqrt{f(r_{ph})}}$ (19)
that $r_{o}$ indicates the observer radius. Shadow radius is shown in Figs. 3 and 4. Table 1, represents the permissible range of parameters based on the observation of Sgr A for $1\sigma$ and $2\sigma$ regions.
Refer to caption      Refer to caption
Figure 3: Variation of shadow radius with parameter $\alpha$ for $r_{o}=10M$. In the left panel $\beta=0.8$ and in the right panel $\Lambda=-0.002$ are considered.
Refer to caption      Refer to caption
Figure 4: Variation of shadow radius with parameter $\Lambda$ for $r_{o}=10M$. In the left panel $\alpha=0.6$ and in the right panel $\beta=0.02$ are considered.
Refer to caption      Refer to caption
Figure 5: Variation of shadow radius with parameter $\beta$ for $r_{o}=10M$. In the left panel $\alpha=0.2$ and in the right panel $\Lambda=-0.002$ are considered.
As evident in Figs. 3 to 5, a reduction in the parameters $\alpha$, and $\Lambda$, while the rest of the parameters remain constant, results in a decrease in the value of the shadow radius.
The allowed range of parameters $\alpha$ and $\Lambda$, where other parameters are constant, in which shadow radius lies in the $1\sigma$ and $2\sigma$ regions, due to the observations of Sgr A is represented in table 1. Also, as shown in Fig. 5, by setting $\alpha=0.2$ and for the given $\Lambda$s, the value of shadow radius lies in the $1\sigma$ region of Sgr A. By setting $\Lambda=-0.002$, for the case $\alpha=0.50$, for $\beta>0.83$, shadow radius is permissible in $1\sigma$ region due to the observations of Sgr A.
Table 1: Upper bounds and lower bounds of the parameters based on the observations of Sgr A.
$1\sigma$ $2\sigma$ $1\sigma$ $2\sigma$
$\alpha\,(\Lambda=-0.002)$ Lower Upper Lower Upper $\alpha\,(\beta=0.800)$ Lower Upper Lower Upper
$\beta=0.900$--- - - ---- - - $0.720$--- - - $\Lambda=-0.002$--- - - ---- - - $0.493$--- - -
$\beta=0.030$--- - - ---- - - $0.508$ $0.813$ $\Lambda=-0.006$--- - - ---- - - $0.959$--- - -
$\beta=0.010$--- - - ---- - - $0.415$ $0.653$ $\Lambda=-0.010$--- - - $0.156$--- - - ---- - -
$\Lambda\,(\beta=0.020)$ Lower Upper Lower Upper $\Lambda\,(\alpha=0.600)$ Lower Upper Lower Upper
$\alpha=0.900$ $-0.041$ $-0.030$ $-0.014$ $-0.007$ $\beta=0.900$ $-0.022$ $-0.014$ $-0.003$--- - -
$\alpha=0.500$ $-0.025$ $-0.017$ $-0.005$--- - - $\beta=0.500$ $-0.024$ $-0.016$ $-0.004$--- - -
$\alpha=0.010$ $-0.015$ $-0.009$--- - - ---- - - $\beta=0.050$ $-0.028$ $-0.019$ $-0.006$ $-0.001$
V Massless Scalar and Electromagnetic Perturbations in Black Hole Spacetimes
In this section, we delve into the dynamics of massless scalar and electromagnetic perturbations within black hole spacetimes. Our analysis operates under the critical assumption that the perturbing fields, whether scalar or electromagnetic, exert a negligible influence on the black hole’s background geometry. This assumption allows us to model the black hole spacetime as a fixed, unperturbed background, focusing instead on the behaviour of the perturbations themselves. This approach is instrumental in understanding the stability of black holes and the nature of the gravitational waves emitted when these astrophysical objects are disturbed.
To investigate these perturbations, we start by deriving Schrödinger-like wave equations of the Klein-Gordon type, which govern the behaviour of massless scalar fields in curved spacetime. These equations are derived from the conservation laws inherent in the black hole’s spacetime, ensuring that the symmetries of the black hole’s geometry are preserved. The resulting wave equations describe how the scalar field evolves as it interacts with the gravitational potential of the black hole, providing a foundational framework for calculating the quasinormal modes (QNMs) associated with these perturbations.
QNMs represent the characteristic oscillations that dominate the response of a black hole to external perturbations. These modes are defined by complex frequencies, which encapsulate both the oscillation frequency and the damping rate of the perturbations. To compute these QNMs, we utilize the higher-order WKB (Wentzel-Kramers-Brillouin) approximation method. The WKB method is a semi-analytical approach that is particularly well-suited for analyzing the complex potentials that arise in black hole spacetimes. It allows for the precise calculation of quasinormal frequencies, especially in the regime where the wavelength of the perturbation is comparable to the size of the black hole’s event horizon.
Understanding the behaviour of QNMs in various black hole spacetimes, including those with modified gravity or additional fields, can enhance our understanding of black holes’ fundamental properties and their role in gravitational wave astronomy. Here we specifically examine massless scalar and vector perturbations within the deformed black hole metric, as described by Eq.(4). Under the assumption that the test fields (whether scalar or vector) have a negligible impact on the black hole’s spacetime, we infer that any back-reaction on the metric is insignificant lopez2020 ; Chandrasekhar:1985kt . To extract the QNMs, we derive Schrödinger-like wave equations for each type of perturbation by applying the relevant conservation laws associated with the spacetime under consideration. The resulting equations are of the Klein-Gordon type for scalar fields and the Maxwell type for electromagnetic fields.
To compute the QNMs, we employ the Padé-averaged 6th-order WKB approximation method. This method provides a suitable framework, allowing for a more comprehensive analysis of the quasinormal spectra associated with different types of perturbations Gogoi:2023kjt ; Gogoi:2024vcx ; Gogoi:2023lvw .
Focusing on axial perturbations, we express the perturbed metric in the form:
$ds^{2}=-g_{tt}dt^{2}+r^{2}\sin^{2}\!\theta\,(d\phi-p_{1}(t,r,\theta)\,dt-p_{2}(t,r,\theta)\,dr-p_{3}(t,r,\theta)\,d\theta)^{2}+g_{rr}\,dr^{2}+r^{2}d\theta^{2},$ (20)
where the functions $p_{1}$, $p_{2}$, and $p_{3}$ represent the perturbations introduced into the black hole spacetime. The metric components $g_{tt}=\mathcal{F}(r)$ and $g_{rr}=1/\mathcal{F}(r)$ correspond to the unperturbed, static, and spherically symmetric components in the perturbative expansion relative to the perturbations $p_{i}$.
By analysing these perturbations within the deformed black hole framework, we aim to further elucidate the stability and dynamical behaviour of black holes under scalar and electromagnetic disturbances. The insights gained from this analysis are expected to contribute to our broader understanding of black hole physics and the role of QNMs in gravitational wave astronomy.
We start by taking into account a massless scalar field in the vicinity of the black hole defined in (6). Since the massless scalar $\Phi$ is assumed to have a negligible back reaction, the stationary Schrödinger equation
$\partial^{2}_{r_{*}}\psi(r_{*})_{sl}+\omega^{2}\psi(r_{*})_{sl}=\mathcal{V}_{s}(r)\psi(r_{*})_{sl},$ (21)
after defining the tortoise coordinate $r_{*}$ as
$\dfrac{dr_{*}}{dr}=\sqrt{g_{rr}g_{tt}^{-1}}=\mathcal{F}(r)$ (22)
such that $\mathcal{V}_{s}(r)$ stands for the effective potential lopez2020
$\mathcal{V}_{s}(r)=|g_{tt}|\left(\dfrac{l(l+1)}{r^{2}}+\dfrac{1}{r\sqrt{|g_{tt}|g_{rr}}}\dfrac{d}{dr}\sqrt{|g_{tt}|g_{rr}^{-1}}\right)=\mathcal{F}(r)\left(\dfrac{l(l+1)}{r^{2}}+\dfrac{1}{r}\dfrac{d}{dr}\mathcal{F}(r)\right)$ (23)
so that, now, $l$ becomes the multipole moment of the black hole’s QNMs. The frequency $\omega$ in the stationary Schrödinger equation (21) is defined via $\partial_{t}^{2}\psi_{sl}(t,r)=-\omega^{2}\psi_{sl}(t,r)$ and this definition is made possible by the diagonal and static nature of the metric in (4). The wave function $\psi(r_{*})_{s}$ stands for the waves associated with the scalar QNMs Chandrasekhar:1985kt ; lopez2020 .
Having done with the scalar perturbations, now we move to the massless vector perturbations, which are nothing but electromagnetic perturbations. In analysing vector perturbations it proves useful to use the tetrad formalism Chandrasekhar:1985kt ; lopez2020 in which the curved metric $g_{\mu\nu}$ in (4) is projected on the flat metric $\eta_{\bar{\mu}\bar{\nu}}$ via the vierbein $e_{\mu}^{\bar{\mu}}$ in the form $g_{\mu\nu}=e_{\mu}^{\bar{\mu}}\eta_{\bar{\mu}\bar{\nu}}e^{\bar{\nu}}_{\nu}$. (Here, the notation is such that while $\mu,\nu,\dots$ refer to curved spacetime, $\bar{\mu},\bar{\nu},\dots$ refer to flat spacetime.) Vector perturbations obey the stationary Schrödinger equation
$\partial^{2}_{r_{*}}\psi_{el}(r_{*})+\omega^{2}\psi_{el}(r_{*})=\mathcal{V}_{em}(r)\psi_{el}(r_{*}),$ (24)
with the effective potential
$\mathcal{V}_{em}(r)=g_{tt}\,\dfrac{l(l+1)}{r^{2}}=\mathcal{F}(r)\dfrac{l(l+1)}{r^{2}}\,.$ (25)
This effective potential $\mathcal{V}_{em}(r)$ of vector perturbations differs from the effective potential $\mathcal{V}_{s}(r)$ of scalar perturbations by the presence of the derivative of $g_{tt}$ in the latter.
Refer to caption      Refer to caption
Figure 6: Behaviour of potentials for scalar and electromagnetic perturbations with $r$.
Refer to caption      Refer to caption
Figure 7: Behaviour of potentials for scalar and electromagnetic perturbations with $r$.
Refer to caption      Refer to caption
Figure 8: Behaviour of potentials for scalar and electromagnetic perturbations with $r$.
In Fig.s 6, 7 and 8, we have depicted the behaviours of the scalar and electromagnetic potentials with respect to the model parameters $l$, $\alpha$ and $\beta$ respectively. From Fig. 6, we can see that the potentials associated with both scalar and electromagnetic potentials behave similarly. However, for the scalar perturbation, the value of the peak of the potential is higher than that of the electromagnetic potential.
Fig. 7 depicts that with an increase in the value of model parameter $\alpha$, the peak value of the potential increases and the peak shifts towards the event horizon of the black hole for both types of perturbations. Finally, in Fig. 8, we have seen that model parameter $\beta$ has an opposite impact on the peak of the potential in comparison to that of the model parameter $\alpha$. So, by investigating the behaviour of the scalar and electromagnetic potentials associated with the black hole spacetime, one can see that both the model parameters $\alpha$ and $\beta$ might have opposite impacts on the QNMs spectrum of the black hole. To understand the impacts of these model parameters on ring-down GWs, we now move to the investigation of QNMs in detail.
Refer to caption      Refer to caption
Figure 9: The variation of QNMs for scalar perturbation with $\alpha$.
Refer to caption      Refer to caption
Figure 10: The variation of QNMs for scalar perturbation with $\beta$.
Refer to caption      Refer to caption
Figure 11: The variation of QNMs for electromagnetic perturbation with $\alpha$.
Refer to caption      Refer to caption
Figure 12: The variation of QNM for electromagnetic perturbation with $\beta$.
This investigation employed the 6th-order WKB approximation technique. Within this method, the oscillation frequency $\omega$ of GWs can be determined using the following expression:
$\omega=\sqrt{-\,i\left[(n+1/2)+\sum_{k=2}^{6}\bar{\Lambda}_{k}\right]\sqrt{-2\mathcal{V}_{0}^{\prime\prime}}+\mathcal{V}_{0}},$ (26)
where $n=0,1,2\ldots$, stands for overtone number, $\mathcal{V}_{0}=\mathcal{V}|_{r\,=\,r_{max}}$ and $\mathcal{V}_{0}^{\prime\prime}=\dfrac{d^{2}\mathcal{V}}{dr^{2}}|_{r\,=\,r_{max}}$. The position $r_{max}$ corresponds to the location where the potential function $\mathcal{V}(r)$ attains its highest value. The correction terms $\bar{\Lambda}_{k}$ play a role in refining the accuracy of the calculations. The specific expressions for these correction terms, along with the Padé averaging method, can be found in the Ref.s Schutz ; Iyer:1986np ; Konoplya:2003ii ; Matyjasek:2019eeu ; Konoplya:2019hlu ; Lambiase:2024lvo ; Konoplya:2023ppx ; Konoplya:2023aph ; Konoplya:2022hll . These references provide detailed information regarding the mathematical forms of the correction terms and the procedure for Padé averaging.
Table 2: The QNMs for the massless scalar and electromagnetic perturbation with $n=0$, $M=1$, $\alpha=0.3$ and $\beta=0.2$ using the Padé averaged 6th order WKB approximation method.
  $l$    Scalar QNMs   Electromagnetic QNMs    $\vartriangle_{rms}(scalar)$    $\vartriangle_{rms}(em)$    $\Delta_{6}(scalar)$    $\Delta_{6}(em)$
$l=1$ $0.310177-0.100055i$ $0.263763-0.0953521i$ $7.68488\times 10^{-7}$ $7.14411\times 10^{-6}$ $3.09809\times 10^{-6}$ $7.38536\times 10^{-6}$
$l=2$ $0.511573-0.0993462i$ $0.484444-0.0977656i$ $2.14225\times 10^{-7}$ $3.75447\times 10^{-7}$ $2.10528\times 10^{-6}$ $5.9346\times 10^{-6}$
$l=3$ $0.714112-0.0991488i$ $0.694861-0.0983548i$ $4.21465\times 10^{-8}$ $5.35672\times 10^{-8}$ $4.10747\times 10^{-7}$ $1.12862\times 10^{-7}$
$l=4$ $0.917035-0.0990674i$ $0.902102-0.0985899i$ $1.06887\times 10^{-8}$ $1.23224\times 10^{-8}$ $1.17441\times 10^{-7}$ $3.25009\times 10^{-7}$
$l=5$ $1.12013-0.0990262i$ $1.10793-0.0987075i$ $3.43245\times 10^{-9}$ $3.78076\times 10^{-9}$ $1.2104\times 10^{-8}$ $1.19967\times 10^{-8}$
We have listed QNMs for $n=0$ and for different $l$ values in Table 2 for both scalar and electromagnetic perturbations. These results were obtained for deformation parameters $\alpha=0.3$ and $\beta=0.2$, which influence the spacetime geometry and black hole properties. The table displays QNM frequencies, with the real part representing the oscillation frequency and the imaginary part corresponding to the damping rate of the modes. These modes were calculated for different multipole numbers $l$, ranging from $l=1$ to $l=5$.
As expected, the real part of the QNMs increases with $l$, indicating higher oscillation frequencies for modes with larger angular momentum. The imaginary part, which controls the damping or decay rate, shows slight variation across the different values of $l$. Interestingly, scalar perturbations consistently exhibit slightly higher oscillation frequencies and more pronounced damping compared to electromagnetic perturbations. This is reflected in the larger imaginary components for scalar modes compared to their electromagnetic counterparts, highlighting the different behaviours of these two types of perturbations in deformed AdS black hole spacetimes.
The table also presents the root mean square (rms) errors, $\Delta_{rms}$, for both scalar and electromagnetic perturbations. These errors quantify the precision of the WKB method in estimating the QNMs. It is observed that the RMS errors generally decrease with increasing $l$, indicating more accurate QNM calculations for higher multipole numbers. The WKB errors, $\Delta_{6}$, which is basically an error term associated with half of the absolute difference between QNMs calculated using 5th and 7th order WKB methods, similarly decrease with increasing $l$. This suggests that the WKB method becomes more reliable for larger values of $l$, where the approximation yields more precise QNM frequencies. Nevertheless, the electromagnetic perturbations exhibit slightly larger errors compared to scalar perturbations, although both remain within acceptable limits, particularly for $l\geq 2$.
In Fig. 9, we have shown the variation of QNMs with respect to the model parameter $\alpha$ for scalar perturbations. On the first panel, we have shown the variations of the real part of QNMs i.e. the oscillation frequencies of ring-down GWs with respect to the model parameter $\alpha$. With an increase in the value of $\alpha$, oscillation frequency increases significantly. In the case of imaginary QNMs, we can see that an increase in the value of $\alpha$ increases the damping rate of ring-down GWs. In Fig. 10, we have depicted the variations of QNMs with respect to the model parameter $\beta$. On the first panel of Fig. 10, one can see that with an increase in the value of $\beta$, the real part of QNMs i.e. oscillation frequency of GWs decreases non-linearly. In the case of damping rate, we observe an interesting behaviour. Initially, the damping rate of the decay rate of ring-down GWs increases with an increase in the value of $\beta$ up to around $\beta=0.65.$ Beyond this value, the decay rate decreases with an increase in the model parameter $\beta$. In the cases of electromagnetic perturbations, we observe a similar scenario with slightly lower values of oscillation frequencies as shown in Fig.s 11 and 12.
The results offer important insights into the physical behaviour of deformed AdS black holes, especially in relation to the parameters $\alpha$ and $\beta$, which modify the spacetime structure and consequently influence the QNM frequencies and damping rates. The distinct responses of the black hole to scalar and electromagnetic perturbations underscore the complexity of black hole interactions with different types of fields. These variations are critical for astrophysical phenomena such as gravitational wave emissions during black hole mergers, where the nature of the perturbing field can impact the ringdown phase. The reliability of the 6th-order WKB method is evident from its accuracy, particularly for higher multipole numbers, affirming its utility in these calculations. The intricate behaviour of the damping rates and oscillation frequencies, particularly the non-linear effects observed with the control parameter $\beta$, suggests that further exploration of these deformation parameters within modified gravity frameworks could deepen our understanding of black hole stability, energy dissipation, and their broader astrophysical and theoretical significance.
VI Greybody Factors
In this section, we examine the calculation of greybody factors in the context of massless scalar and electromagnetic perturbations in black hole spacetimes. Building on Hawking’s groundbreaking discovery in 1975, which revealed that black holes emit radiation known as Hawking radiation, subsequent research has demonstrated that this radiation is not completely black but rather has a greybody appearance due to modifications in the spectrum as it travels away from the black hole’s horizon Hawking:1975vcx ; Singleton:2011vh ; Akhmedova:2008dz . This phenomenon, characterized by the greybody factor, accounts for the redshift and other effects that alter the observed radiation compared to its original form near the event horizon.
To calculate these greybody factors, various methods have been developed and utilized in the literature, such as those by Maldacena et al. (1996) and Fernando (2004), among others Maldacena:1996ix ; Fernando:2004ay ; Okyay2022 ; Ovgun188 ; Pantig:2022gih ; Yang:2022xxh ; Panotopoulos:2018pvu ; Panotopoulos:2016wuu ; Rincon:2018ktz ; Ahmed:2016lou ; Javed:2022kzf ; Al-Badawi:2022aby . In this study, we employ the higher-order WKB (Wentzel-Kramers-Brillouin) approximation method to determine greybody factors for scalar and electromagnetic perturbations.
We begin by analyzing the wave equation (21) under boundary conditions that allow for incoming waves from infinity. Given the symmetry of scattering properties, this is equivalent to considering the scattering of a wave originating from the black hole horizon. This setup is particularly useful for determining the proportion of waves reflected back towards the horizon by the potential barrier. The boundary conditions for this scenario are expressed as:
$\begin{array}[]{ccll}\Psi&=&e^{-i\omega r_{*}}+Re^{i\omega r_{*}},&r_{*}\rightarrow+\infty,\\ \Psi&=&Te^{-i\omega r_{*}},&r_{*}\rightarrow-\infty,\\ \end{array}$ (27)
where $R$ and $T$ denote the reflection and transmission coefficients, respectively, and satisfy the conservation of probability:
$\left|T\right|^{2}+\left|R\right|^{2}=1.$ (28)
The transmission coefficient $\left|T\right|$, which is equivalent to the greybody factor $A$ of the black hole, is calculated using the WKB approach:
$\left|A\right|^{2}=1-\left|R\right|^{2}=\left|T\right|^{2}.$ (29)
In the WKB method, the reflection coefficient $R$ is given by:
$R=(1+e^{-2i\pi K})^{-\frac{1}{2}},$ (30)
where the phase factor $K$ is determined by the equation:
$K-i\frac{(\omega^{2}-\mathcal{V}_{0})}{\sqrt{-2\mathcal{V}_{0}^{\prime\prime}}}-\sum_{i=2}^{i=6}\Lambda_{i}(K)=0.$ (31)
Here, $\mathcal{V}_{0}$ represents the maximum value of the effective potential, $\mathcal{V}_{0}^{\prime\prime}$ is its second derivative at the maximum, and $\Lambda_{i}$ denotes higher-order WKB corrections that depend on the potential’s derivatives up to the $2i$th order at the maximum. These corrections, as discussed in works such as Schutz ; Will_wkb ; Konoplya_wkb ; Maty_wkb , are crucial for improving the accuracy of the WKB approximation. Although the 6th order WKB formula is the primary method used in this study, lower orders may be applied for lower frequencies and multipole numbers.
It is important to note that the WKB approach’s accuracy diminishes at low frequencies, where the reflection is nearly total and the greybody factors approach zero. However, this limitation does not significantly affect the estimation of energy emission rates, which remains accurate under the WKB framework. Since the WKB method is a well-established technique, further details are omitted here; for more information, readers are referred to comprehensive reviews such as Konoplya:2019hlu ; Konoplya:2011qq .
Refer to caption      Refer to caption
Figure 13: The greybody for scalar and electromagnetic perturbation with $\omega$
Refer to caption      Refer to caption
Figure 14: The variation of greybody for scalar and electromagnetic perturbation with $\omega$
Refer to caption      Refer to caption
Figure 15: The variation of greybody for scalar and electromagnetic perturbation with $\omega$
We have shown the variations of greybody factors with respect to the model parameters in Fig.s 13, 14 and 15.
These figures illustrate the behaviour of greybody factors for scalar and electromagnetic (EM) perturbations in deformed black holes. The plots explore the effects of various parameters, including the deformation parameter $\alpha$, control parameter $\beta$, and the angular momentum quantum number $l$, on the absorption probability $|A|^{2}$ as a function of frequency $\omega$. These figures shed light on how these parameters influence the transmission of perturbations through the potential barrier surrounding the black hole, offering insights into the black hole’s interaction with its surroundings.
In Fig. 13, the greybody factors are examined for different values of $l$ and $\alpha$, with $\beta$ held constant. The two panels depict the behaviour of scalar and electromagnetic perturbations for $l=1$ to $l=4$, showing that higher angular momentum perturbations face a higher potential barrier, leading to a steeper ascent in the greybody factor curves at higher frequencies. This suggests that perturbations with greater angular momentum are more strongly scattered by the potential, reducing their likelihood of transmission at lower frequencies.
The two panels of Fig. 14 focus on the effect of varying $\alpha$, with $l=2$ and $\beta=0.2$ held constant. As $\alpha$ increases, the onset of significant transmission shifts to higher frequencies, indicating that the deformation of the black hole spacetime increases the potential barrier. This behaviour highlights the strong influence of the deformation parameter on the greybody factors, suggesting that modifications to black hole geometry can have a noticeable effect on the transmission characteristics of perturbations.
Fig. 15 further explores the impact of the control parameter $\beta$, with fixed values for $\alpha=0.3$ and $l=2$. The greybody factors for scalar and electromagnetic perturbations are plotted for $\beta$ values ranging from 0.1 to 1.3. Here, the influence of $\beta$ is more subtle compared to $\alpha$ or $l$, as the curves for different $\beta$ values are closely aligned. While increasing $\beta$ leads to a slightly earlier onset of transmission and a marginally sharper rise in $|A|^{2}$, the overall effect of $\beta$ on the greybody factor is minor in this frequency range.
When combining the results from all these three figures, it is clear that the angular momentum quantum number $l$ and the deformation parameter $\alpha$ play a more prominent role in shaping the greybody factors, significantly affecting the absorption probabilities of perturbations. The control parameter $\beta$, while still influential, fine-tunes the greybody factor profiles without causing dramatic shifts. This suggests that the most notable observational signatures of deformed black holes might arise from variations in $l$ and $\alpha$, while changes in $\beta$ would have more nuanced effects.
Overall, these results provide a comprehensive view of how greybody factors depend on key parameters in deformed black hole spacetimes. The findings have important implications for understanding the observational signatures of black holes and the influence of modified gravity theories on black hole properties. Further investigations into the interplay of these parameters could help refine models of black hole evaporation, quasinormal modes, and related phenomena, contributing to a deeper understanding of gravity in extreme regimes.
VII Emission Rate
The particle emission rate from a black hole is a critical optical property that provides insight into the black hole’s stability and lifetime. This emission rate is intrinsically linked to two fundamental characteristics: the Hawking temperature and the black hole shadow. The Hawking temperature, which arises from quantum effects near the event horizon, governs the rate at which particles are thermally emitted by the black hole. Meanwhile, the shadow - the dark region against the backdrop of luminous sources - provides a geometric measure that influences the observational properties of black hole emissions. Together, these attributes enable a comprehensive understanding of the energy emission profile, which can be expressed as Wei2013 ; Cai:2020igv ; Decanini:2011xi
$\frac{\partial^{2}E}{\partial t\partial\omega}=\frac{\left(2\pi^{2}\omega^{3}\right)\sigma}{e^{\frac{\omega}{T_{H}}}-1}$ (32)
where $\sigma=\pi r_{sh}^{2}$ represents the cross section, $r_{sh}$ and $T_{H}$ are obtained from Eqs. (19) and (8), respectively. Analyzing this emission rate allows researchers to evaluate the black hole’s dynamic properties and to predict its longevity, with significant implications for studying black hole thermodynamics and quantum gravity.
Refer to caption      Refer to caption
Figure 16: Variation of emission rate. Left panel: Parameters are set as $M=1,\,\beta=0.2,$ and $\Lambda=-0.002$. Right panel: Parameters are set as $M=1,\,\alpha=0.7,$ and $\Lambda=-0.002$
As shown in Fig. 16, an increase in the parameter $\alpha$, causes the maximum value of the emission rate to decrease and occur at a lower frequency. Also, an increase in the parameter $\beta$, leads to a decrease in the emission rate maximum value and causes the maximum occurring at a higher frequency value.
VIII Concluding Remarks
In this work, we explored several properties of the deformed black hole spacetime such as the temperature, photon and shadow radii, QNMs, greybody factors and emission rate, mainly focusing on the effects of the deformation parameter $\alpha$ and control parameter $\beta$ on both scalar and electromagnetic perturbations. The investigation of shadow radius as a function of parameters $\alpha,\,\beta$ and $\Lambda$, leads to an upper and lower limit of parameters in which the shadow radius is allowed as a result of the observed data of Sgr A. Our analysis of QNMs revealed that an increase in $\alpha$ significantly enhances the oscillation frequency and damping rate of ringdown GWs while increasing $\beta$ leads to a non-linear behaviour in both frequency and damping rates. Specifically, the damping rate initially increases with $\beta$ before reversing its trend beyond $\beta=0.65$. For electromagnetic perturbations, we observed a similar pattern, though with slightly lower oscillation frequencies and damping rates compared to scalar perturbations. These variations underscore the sensitivity of black hole dynamics to these parameters, which are crucial for understanding astrophysical processes like black hole mergers and gravitational wave emissions.
Additionally, we examined the greybody factors for scalar and electromagnetic perturbations, with a focus on the role of angular momentum $l$, $\alpha$, and $\beta$. Our results indicate that both $l$ and $\alpha$ play a dominant role in shaping the transmission of perturbations through the black hole’s potential barrier. Higher angular momentum leads to stronger scattering while increasing $\alpha$ shifts the onset of significant transmission to higher frequencies. In contrast, $\beta$ has a more subtle impact, fine-tuning the transmission profiles without causing major shifts in the greybody factors.
Overall, our findings highlight the intricate interplay between deformation parameters and the behaviour of black holes under perturbations. The results provide valuable insights into how modified black hole geometries affect QNMs, damping rates, and greybody factors, with implications for both theoretical studies of black holes and their observational signatures in gravitational wave astronomy. Future work could focus on exploring these effects within different modified gravity frameworks to further our understanding of black hole evolution and energy dissipation in extreme regimes.
Acknowledgments
DJG acknowledges the contribution of the COST Action CA21136 – “Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse)”. Also, the research was partially supported by the Long-Term Conceptual Development of a University of Hradec Králové for 2023, issued by the Ministry of Education, Youth, and Sports of the Czech Republic.
Data Availability Statement
There are no new data associated with this article.